# BIOSCAN-5M: A Multimodal Dataset for Insect Biodiversity

### v1

@@ Acknowledgement
1 | processid | An identifier given by the collector. | String |
2 | sampleid | An identifier given by the collector. | String |
3 | taxon | Bio.info: Most specific taxonomy rank. | String |
4 | phylum | Bio.info: Taxonomic classification label at phylum rank. | String |
5 | class | Bio.info: Taxonomic classification label at class rank. | String |
6 | order | Bio.info: Taxonomic classification label at order rank. | String |
7 | family | Bio.info: Taxonomic classification label at family rank. | String |
8 | subfamily | Bio.info: Taxonomic classification label at subfamily rank. | String |
9 | genus | Bio.info: Taxonomic classification label at genus rank. | String |
10 | species | Bio.info: Taxonomic classification label at species rank. | String |
11 | dna_bin * | Bio.info: Barcode Index Number (BIN). | String |
12 | dna_barcode * | Bio.info: Nucleotide barcode sequence. | String |
13 | country | Geo.info: Country associated with the site of collection. | String |
14 | province_state | Geo.info: Province/state associated with the site of collection. | String |
15 | coord-lat | Geo.info: Latitude (WGS 84; decimal degrees) of the collection site. | Float |
16 | coord-lon | Geo.info: Longitude (WGS 84; decimal degrees) of the collection site. | Float |
17 | image_measurement_value | Size.info: Number of pixels occupied by the organism. | Integer |
18 | area_fraction | Size.info: Fraction of the original image the cropped image comprises. | Float |
19 | scale_factor | Size.info: Ratio of the cropped image to the cropped_256 image. | Float |
20 | inferred_ranks | An integer indicating at which taxonomic ranks the label is inferred. | Integer |
21 | split | Split set (partition) the sample belongs to. | String |
22 | index_bioscan_1M_insect | An index to locate organism in BIOSCAN-1M Insect metadata. | Integer |
23 | chunk | The packaging subdirectory name (or empty string) for this image. | String |
@@
microMalaise01 | microMalaise01 Malaise1237 | 13 |
@@
We pretrain the model on the 2,283,900 unique DNA sequences from the pretrained partition and the 41,232 unique sequences from the other_heldout partition, totalling 2,325,132 pretraining DNA samples. For all samples, trailing N
characters are removed and all sequences are truncated at 660 nucleotides. Note that leading N
characters are retained since they are likely to correspond to true unknown nucleotides in the barcode. The model was pretrained using the same MLM loss function and training configurations as in BarcodeBERT [52]. Specifically, we use a non-overlapping -mer-based tokenizer and a transformer model with 12 transformer layers, each having 12 attention heads. However, we included a random offset of at most nucleotides to each sequence as a data augmentation technique to enhance the sample efficiency. We use a learning rate of , a batch size of 128, a OneCycle scheduler [67], and the AdamW optimizer [46], training the model for 35 epochs. In addition to using the architecture reported in BarcodeBERT, we performed a parameter search to determine the optimal -mer tokenization length and model size, parameterized by the number of layers and heads in the transformer model, in order to identify an optimal architecture configuration. After pretraining, we fine-tuned the model with cross-entropy supervision for species-level classification. The pre-training stage takes approximately 50 hours using four Nvidia A40 GPUs and the fine-tuning stage of the 4-12-12 models takes 2.5 hours in four Nvidia A40 GPUs.
@@
4 layers, 4 heads | 6 layers, 6 heads | 12 layers, 12 heads | ||||||||||

### v3

@@ Acknowledgement
| 1 | processid | An identifier given by the collector. | String |
| 2 | sampleid | An identifier given by the collector. | String |
| 3 | taxon | Bio.info: Most specific taxonomy rank. | String |
| 4 | phylum | Bio.info: Taxonomic classification label at phylum rank. | String |
| 5 | class | Bio.info: Taxonomic classification label at class rank. | String |
| 6 | order | Bio.info: Taxonomic classification label at order rank. | String |
| 7 | family | Bio.info: Taxonomic classification label at family rank. | String |
| 8 | subfamily | Bio.info: Taxonomic classification label at subfamily rank. | String |
| 9 | genus | Bio.info: Taxonomic classification label at genus rank. | String |
| 10 | species | Bio.info: Taxonomic classification label at species rank. | String |
| 11 | dna_bin * | Bio.info: Barcode Index Number (BIN). | String |
| 12 | dna_barcode * | Bio.info: Nucleotide barcode sequence. | String |
| 13 | country | Geo.info: Country associated with the site of collection. | String |
| 14 | province_state | Geo.info: Province/state associated with the site of collection. | String |
| 15 | coord-lat | Geo.info: Latitude (WGS 84; decimal degrees) of the collection site. | Float |
| 16 | coord-lon | Geo.info: Longitude (WGS 84; decimal degrees) of the collection site. | Float |
| 17 | image_measurement_value | Size.info: Number of pixels occupied by the organism. | Integer |
| 18 | area_fraction | Size.info: Fraction of the original image the cropped image comprises. | Float |
| 19 | scale_factor | Size.info: Ratio of the cropped image to the cropped_256 image. | Float |
| 20 | inferred_ranks | An integer indicating at which taxonomic ranks the label is inferred. | Integer |
| 21 | split | Split set (partition) the sample belongs to. | String |
| 22 | index_bioscan_1M_insect | An index to locate organism in BIOSCAN-1M Insect metadata. | Integer |
| 23 | chunk | The packaging subdirectory name (or empty string) for this image. | String |
@@
| microMalaise01 | microMalaise01 Malaise1237 | 13 | ||||
@@
We pretrain the model on the 2,283,900 unique DNA sequences from the pretrained partition and the 41,232 unique sequences from the other_heldout partition, totalling 2,325,132 pretraining DNA samples. For all samples, trailing N characters are removed and all sequences are truncated at 660 nucleotides. Note that leading N characters are retained since they are likely to correspond to true unknown nucleotides in the barcode. The model was pretrained using the same MLM loss function and training configurations as in BarcodeBERT [52]. Specifically, we use a non-overlapping -mer-based tokenizer and a transformer model with 12 transformer layers, each having 12 attention heads. However, we included a random offset of at most nucleotides to each sequence as a data augmentation technique to enhance the sample efficiency. We use a learning rate of , a batch size of 128, a OneCycle scheduler [67], and the AdamW optimizer [46], training the model for 35 epochs. In addition to using the architecture reported in BarcodeBERT, we performed a parameter search to determine the optimal -mer tokenization length and model size, parameterized by the number of layers and heads in the transformer model, in order to identify an optimal architecture configuration. After pretraining, we fine-tuned the model with cross-entropy supervision for species-level classification. The pre-training stage takes approximately 50 hours using four Nvidia A40 GPUs and the fine-tuning stage of the 4-12-12 models takes 2.5 hours in four Nvidia A40 GPUs.
@@
| 4 layers, 4 heads | 6 layers, 6 heads | 12 layers, 12 heads | ||||||||||